OpenStudy (anonymous):

f(x) ln|x^2-11| Differentiate. (derivatives of logarithmic functions)

OpenStudy (anonymous):

\[f'(x)=\frac{ 1 }{ x ^{2}-11 }2x\]

OpenStudy (anonymous):

=(2 x)/(-11+x^2)

OpenStudy (anonymous):

sorry guys there was a typo. not eleven but 1.

OpenStudy (anonymous):

can u show me the steps?

OpenStudy (anonymous):

do you know the chain rule ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what it says ?

OpenStudy (anonymous):

are u asking what the chain rule is? not sure what you mean ...

OpenStudy (anonymous):

chain rule in differentiation

OpenStudy (anonymous):

if we want to differentiate a function (say f(x)), in your case it is f(x)= ln|x^2-11|. What will be the first step ?

OpenStudy (anonymous):

i actually don't have the chain rule quite down. could you show me? also its number one instead of eleven. it was a typo

OpenStudy (anonymous):

let f(x)= ln|g(x)| what did you get differentiating this

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

nopes. f'(x)={1/g(x)} g'(x) put g(x)= x^2-1

OpenStudy (anonymous):

then what will be the g'(x) ?

OpenStudy (anonymous):

1/2x * 2x

OpenStudy (anonymous):

wouldn't g'x just be 2x because x^2 = 2x and -1 = 0?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but our f'(x) is \[f(x)=\frac{ 1 }{ g(x) }g'(x)\]

OpenStudy (anonymous):

@sohrinny Do you still need help or are you clear on what is to be done?

OpenStudy (anonymous):

no i still need help